OpenStudy (anonymous):

solve the non standard equation for x, 17=9+x .. ? help with this problem ?? .....,.....

OpenStudy (anonymous):

Subtract 9 from both sides of the equation.

OpenStudy (anonymous):

-9 + 17=x so, x=8 :)

OpenStudy (anonymous):

\[17=9+x \implies 17-9=9-9+x \implies 8=x\]